A) do e and m vary directly,inversely,or neither?
e)/1/2/3
m)/36/18/12 (this is a table)

B) if e and m vary directly or inversely,write a formula to express the relationship between them. if there is no relationship,explain why.

Answer by Alan3354(69443)

A) do e and m vary directly,inversely,or neither?
e)/1/2/3
m)/36/18/12 (this is a table)
--------------
e*m = 36 for the 3 points given.
--> e = 36/m an inverse relationship.
--------------------------
B) if e and m vary directly or inversely,write a formula to express the relationship between them.
I did.
e = 36/m or m = 36/e
---------------------
if there is no relationship,explain why.